OpenStudy (anonymous):

Identify the 12th term of a geometric sequence where a1 = 8 and a6 = –8,192.

OpenStudy (anonymous):

do you have a2

OpenStudy (anonymous):

no i dont

OpenStudy (amistre64):

since youve gone thru this with alot of people already ... what is the general equation for an arith. progression ?

OpenStudy (anonymous):

a2=-32

OpenStudy (anonymous):

an= ar1 * r^(n-1) this is the geometric equation

OpenStudy (amistre64):

lol ... its just too easy to read over that word :) yes, this is a geo ... good eye

OpenStudy (anonymous):

lol, i see it now

OpenStudy (anonymous):

a12=-33554432

OpenStudy (anonymous):

this is how i found it, give me a moment i will explain

OpenStudy (anonymous):

since the formula for any term is given by an= a1 *r^(n-1) in order to find any term you must know a1 and r you have a1 but we need to find r, in order to find r we need an and a1 we only have a1 and a6 so substitute a6 for an in the above formula a6=8*r ^(6-1) -8192=8 * r ^(5) -1024=r^5 -4=r now using above formula again we have a12=8 * (-4)^(12-1) a12=8 * (-4)^11 a12=8 * (-4194304) a12=-33554432 hint* any time a term is negative, the ratio will be negative

OpenStudy (anonymous):

formula for finding any term in ageometric sequence an=a1*r^(n-1) i have never worked a related problem, so that is why it took so long, but now we both know amistre64 helped me realize you needed a12 good day

OpenStudy (anonymous):

geometric equation a.k.a arithmetic transformation